What is the quadratic equation of 4x^2-8x-9=0
(simplifiy your answer. type an exact answer using radicals as needed. Use integers or fractions for any number in the expression. Express complex numbers in the form of i.)

Answer by mathitutor(25)

4x^2-8x-9=0
=> (2x)^2 - 2.2x.2 + 2^2 -9 -4 = 0
=> (2x -2)^2 = 13
=> 2x - 2 = +sqrt(13) or -sqrt(13)
=> 2x = 2 +sqrt(13) or 2-sqrt(13)
x = 1 + sqrt(13)/2 or 1 - sqrt(13)/2